OpenStudy (anonymous):

If h(x)=(f o g)(x) and h(x)= 4sqrt x+7, find g(x) if f(x)= 4sqrt x+1 plz help. I dont know how to start it

OpenStudy (anonymous):

\[\text {Is }h(x) = 4\sqrt{x + 7} \text{and }f(x) = 4\sqrt{x + 1}?\]

OpenStudy (anonymous):

Is that what you meant?

OpenStudy (anonymous):

no hold ona sec

OpenStudy (anonymous):

Ok. h(x) = f(g(x)) \[\sqrt[4]{x + 7} = \sqrt[4]{g(x) + 1}\]\[(\sqrt[4]{x + 7})^{4} = (\sqrt[4]{g(x) + 1})^{4}\]\[x + 7 = g(x) + 1\]\[g(x) = x + 6\]

OpenStudy (anonymous):

so u simplfy them an then...

OpenStudy (anonymous):

That's it. That's what g(x) is.

OpenStudy (anonymous):

so now u do \[\sqrt[4]{x+1}\]*x=6?

OpenStudy (anonymous):

Reread what the problem says. It doesn't say to solve for x.

OpenStudy (anonymous):

oh nooo solve for g(x) and that x+6 sorry -_-

OpenStudy (anonymous):

lol. It's ok :)

OpenStudy (anonymous):

so x+6 is the answer

OpenStudy (anonymous):

Yup. g(x) = x + 6

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

np :)